OpenStudy (anonymous):

Find the zeros of the function h(x) = x2 +4x by factoring. What is the smallest zero of the function? Enter a number only.

OpenStudy (agreene):

x(x+4)

OpenStudy (anonymous):

-4

OpenStudy (agreene):

x=-4 and x=0

OpenStudy (ash2326):

We have \[h(x)=x^2+4x\] We have to find values of x for which h(x) is zero Let' put h(x)=0 \[0= x^2+4x\] \[x^2+4x=0\] \[x(x+4)=0\] so the zeroes are x=0 and x=-4 The smaller one is your answer